OpenStudy (anonymous):

Find three consecutive even intergers whose sum is 114.

OpenStudy (anonymous):

i need it to be in a linear equation please.

OpenStudy (anonymous):

oh wait even

OpenStudy (anonymous):

x+x+2+x+4=114

OpenStudy (anonymous):

x+x+2+x+4=114

OpenStudy (anonymous):

how would i solve that? :o

OpenStudy (anonymous):

add like terms

OpenStudy (anonymous):

3x+6=114

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):